### v1

# exampleloeList of Examples

Abstract.
Designing preference elicitation (PE) methodologies that can quickly ascertain a user’s top item preferences in a cold-start setting is a key challenge for building effective and personalized conversational recommendation (ConvRec) systems. While large language models (LLMs) constitute a novel technology that enables fully natural language (NL) PE dialogues, we hypothesize that monolithic LLM NL-PE approaches lack the multi-turn, decision-theoretic reasoning required to effectively balance the NL exploration and exploitation of user preferences towards an arbitrary item set. In contrast, traditional Bayesian optimization PE methods define theoretically optimal PE strategies, but fail to use NL item descriptions or generate NL queries, unrealistically assuming users can express preferences with direct item ratings and comparisons. To overcome the limitations of both approaches, we formulate NL-PE in a Bayesian Optimization (BO) framework that seeks to actively elicit natural language feedback to reduce uncertainty over item utilities to identify the best recommendation. Key challenges in generalizing BO to deal with natural language feedback include determining (a) how to leverage LLMs to model the likelihood of NL preference feedback as a function of item utilities and (b) how to design an acquisition function that works for language-based BO that can elicit in the infinite space of language. We demonstrate our framework in a novel NL-PE algorithm, PEBOL, which uses Natural Language Inference (NLI) between user preference utterances and NL item descriptions to maintain preference beliefs and BO strategies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) to guide LLM query generation. We numerically evaluate our methods in controlled experiments, finding that PEBOL achieves up to 131% improvement in MAP@10 after 10 turns of cold start NL-PE dialogue compared to monolithic GPT-3.5, despite relying on a much smaller 400M parameter NLI model for preference inference.
Conversational Recommendation, Preference Elicitation, Bayesian Optimization, Online Recommendation, Query Generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Personalizationccs: Information systems Language models
1. Introduction
Personalized conversational recommendation (ConvRec) systems require effective natural language (NL) preference elicitation (PE) strategies that can efficiently learn a user’s top item preferences in cold start settings, ideally requiring only an arbitrary set of NL item descriptions. While the advent of large language models (LLMs) has introduced the technology to facilitate NL-PE conversations (Handa et al., 2024; Li et al., 2023) we conjecture that monolithic LLMs have limited abilities to strategically conduct active, multi-turn NL-PE dialogues about a set of arbitrary items. Specifically, we hypothesize that LLMs lack the sequential decision-theoretic reasoning to interactively generate queries that avoid over-exploitation or over-exploration of user-item preferences, thus risking over-focusing on already revealed item preferences or wastefully exploring preferences over low-value items. Further challenges faced by monolithic LLM NL-PE approaches include the need to jointly reason over large, potentially unseen sets of item descriptions, and the lack of control and interpretability in system behaviour even after prompt engineering or fine-tuning (Maynez et al., 2020).
Figure 1. PEBOL’s belief updates over a cold-start user’s item utilities during three turns of NL dialogue. Bayesian preference beliefs not only facilitate recommendation, but also enable Bayesian optimization policies to guide LLM query generation, avoiding over-exploration (asking about clearly low-value items) and over-exploitation (over-focusing on known preferences).
In contrast, conventional PE algorithms (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020; Lee et al., 2019; Zhao et al., 2013), including Bayesian optimization methods (Guo and Sanner, 2010; Yang et al., 2021; Christakopoulou et al., 2016; Vendrov et al., 2019; Boutilier, 2002), establish formal decision-theoretic policies such as Thompson Sampling (TS) and Upper Confidence Bound (UCB) (Karamanolakis et al., 2018) to balance exploration and exploitation with the goal of quickly identifying the user’s most preferred items. However, these techniques typically assume a user can express preferences via direct item ratings or comparisons – an unrealistic expectation when users are unfamiliar with most items (Biyik et al., 2023). While recent work has extended Bayesian PE to a fixed set of template-based queries over pre-defined keyphrases (Yang et al., 2021), no existing work extends Bayesian methodologies to generative NL-PE over a set of generic NL item descriptions.
In this paper, we make the following contributions:
  • We introduce the first Bayesian optimization formalization of NL-PE for arbitrary NL dialogue over a generic set of NL item descriptions – establishing a framework for research on augmenting LLMs with decision-theoretic reasoning.
  • We present PEBOL (Preference Elicitation with Bayesian Optimization augmented LLMs), a novel NL-PE algorithm which 1) infers item preferences via Natural Language Inference (NLI) (Yin et al., 2019) between dialogue utterances and item descriptions to maintain Bayesian preference beliefs and 2) introduces LLM-based acquisition functions, where NL query generation is guided by decision-theoretic strategies such as TS and UCB over the preference beliefs.
  • We numerically evaluate PEBOL against monolithic GPT-3.5 NL-PE methods via controlled NL-PE dialogue experiments over multiple NL item datasets and levels of user noise.
  • We observe that PEBOL achieves up to 131% improvement in MAP@10 after 10 turns compared to GPT-3.5, despite using a much smaller 400M NLI model for preference inference.
2. Background and Related Work
2.1. Bayesian Optimization
Given an objective function $f:\mathcal{X}\rightarrow\mathbb{R}$, (standard) optimization systematically searches for a point $x^{*}\in\mathcal{X}$ that maximizes [^1] We take the maximization direction since this paper searches for items with maximum utility for a person. $f$. Bayesian optimization focuses on settings where $f$ is a black-box function which does not provide gradient information and cannot be evaluated exactly – rather, $f$ must be evaluated using indirect or noisy observations which are expensive to obtain (Garnett, 2023; Shahriari et al., 2016). To address these challenges, Bayesian optimization maintains probabilistic beliefs over $f(x)$ and its observations to guide an uncertainty-aware optimization policy which decides where to next observe $f(x)$.
Bayesian optimization begins with a prior $p(f)$ which represents the beliefs about $f$ before any observations are made. Letting $y_{i}$ represent a noisy or indirect observation of $f(x_{i})$, and collecting a sequence of observations into a dataset $\mathcal{D}=(\mathbf{x},\mathbf{y})$, an observation model defines the likelihood $p(\mathcal{D}|f)$. We then use the observed data and Bayes theorem to update our beliefs and obtain the posterior
(1) $p(f|\mathcal{D})=\frac{p(f)p(\mathcal{D}|f)}{p(\mathcal{D})}.$
This posterior informs an acquisition function $\gamma(x|\mathcal{D})$ which determines where to next observe $f(x)$ in a way that balances exploitation (focusing observations where $f$ is likely near its maximum) with exploration (probing areas where $f$ has high uncertainty).
2.2. Preference Elicitation
PE has witnessed decades of research, and includes approaches based on Bayesian optimization (e.g., (Eric et al., 2007; Guo et al., 2010; Brochu et al., 2010; González et al., 2017; Khajah et al., 2016)), Bandits (e.g., (Li et al., 2010, 2011; Zhao et al., 2013; Christakopoulou et al., 2016)), constrained optimization (Rossi and Sperduti, 2004), and POMDPs (Boutilier, 2002). In the standard PE setting, a user is assumed to have some hidden utilities $\mathbf{u}=[u_{1},...,u_{N}]$ over a set $\mathcal{I}$ of $N$ items, where item $i$ is preferred to item $j$ if $u_{i}>u_{j}$. The goal of PE is typically to search for an item $i^{*}\in\arg\max_{i}u_{i}$ that maximizes user utility in a minimal number of PE queries, which most often ask a user to express item preferences as item ratings (e.g., (Li et al., 2010, 2011; Brochu et al., 2010; Zhao et al., 2013; Christakopoulou et al., 2016)) or relative preferences between item pairs or sets (e.g., (Boutilier, 2002; Guo and Sanner, 2010; Handa et al., 2024; Eric et al., 2007; González et al., 2017; Vendrov et al., 2019)). An alternative form of PE asks users to express preferences over predefined item features, also through rating- or comparison-based queries (Li et al., 2021; Zhang et al., 2020; Lei et al., 2020). Central to the above PE methods are query selection strategies that balance the exploration and exploitation of user preferences, with TS and UCB algorithms (cf. Sec. 4.2) often exhibiting strong performance (Zhao et al., 2013; Christakopoulou et al., 2016; Yang et al., 2021; Li et al., 2021; Zhang et al., 2020). However, none of these methods are able to interact with users through NL dialogue.
Figure 2. The PEBOL NL-PE algorithm, which maintains a Bayesian belief state over a user’s item preferences given an arbitrary set of NL item descriptions $\mathbf{x}$. This belief is used by a decision-theoretic policy to balance the exploration and exploitation of preferences by strategically selecting an item description $x_{i^{t}}$ as the basis for LLM query generation. Belief updates are computed through Bayesian inference with NLI entailment scores between item descriptions and query-response pairs.
2.3. Language-Based Preference Elicitation
Yang et al. (Yang et al., 2021) introduce Bayesian PE strategies using TS and UCB for keyphrase rating queries, where keyphrases are first mined from NL item reviews and then co-embedded with user-item preferences in a recommendation system. Handa et al. (Handa et al., 2024) propose using LLMs to interface with a conventional Bayesian PE system, suggesting a preprocessing step to extract features from NL descriptions and a verbalization step to fluidly express pairwise item comparison queries. Li et al. (Li et al., 2023) prompt an LLM to generate PE queries for some specific domain (e.g., news content, morals), observe user responses, and evaluate LLM relevance predictions for a single item. While these works make progress towards NL-PE, they do not study how LLM query generation can strategically explore user preferences towards an arbitrary item set outside the realm of item-based or category-based feedback.
2.4. Conversational Recommendation
Recent work on ConvRec uses language models [^2] Earlier systems (e.g. (Li et al., 2018; Chen et al., 2019)) use relatively small RNN-based language models. to facilitate NL dialogue while integrating calls to a recommender module which generates item recommendations based on user-item interaction history (Li et al., 2018; Chen et al., 2019; Wang et al., 2022; Yang et al., 2022). He et al. (He et al., 2023) report that on common datasets, zero-shot GPT-3.5/4 outperforms these ConvRec methods, which generally use older language models and require user-item interaction history for their recommendation modules.
2.5. Natural Language Inference
Binary Natural Language Inference (NLI) (Yin et al., 2019) models predict the likelihood that one span of text called a premise is entailed by (i.e., can be inferred from) a second span called the hypothesis. For example, an effective NLI model should predict a high likelihood that the premise “I want to watch Iron Man” entails the hypothesis “I want to watch a superhero movie”. As illustrated by this example, the hypothesis typically must be more general than the premise. NLI models are trained by fine-tuning encoder-only LLMs on NLI datasets (Williams et al., 2017; Dagan et al., 2005; Thorne et al., 2018), which typically consist of short text spans for the premise and hypothesis – thus enabling relatively efficient performance on similar tasks with a fairly small number LLM parameters.
Figure 3. Cherry-picked system-generated dialogues from our NL-PE experiments. The Monolithic GPT-3.5 dialogue (left) demonstrates over-exploitation, with $q^{3}$ directly extending $q^{2}$ after a positive user preference is observed and leading to the extreme case of query repetition ($q^{4}$ = $q^{3}$). In contrast, PEBOL (right) continues exploring even after a positive response, while focusing on promising aspects (three out of four queries elicit a positive response) by using UCB-guided query generation.
3. Problem Definition
We now present a Bayesian optimization formulation of NL-PE. The goal of NL-PE is to facilitate a NL dialogue which efficiently discovers a user’s most preferred items out of a set of $N$ items. Each item $i\in\mathcal{I}$ has a NL description $x_{i}$, which might be a title, long-form description, or even a sequence of reviews, with the item set $\mathcal{I}$ collectively represented by $\mathbf{x}\in\mathcal{X}$ with $\mathbf{x}=[x_{1},...,x_{N}]$. We assume the user has some (unknown) utility function $f:\mathcal{X}\rightarrow\mathbb{R}$ establishing hidden utilities $\mathbf{u}=f(\mathbf{x})$ so that item $i$ is preferred to item $j$ if $u_{i}>u_{j}$. Our goal is to find the most preferred item(s):
(2) $i^{*}\in\arg\max_{i\in\mathcal{I}}u_{i}.$
In contrast to standard Bayesian PE formalisms (c.f. Sec 2.2), we do not assume that the user can effectively convey direct item-level preferences by either: 1) providing item ratings (i.e., utilities) or 2) pairwise or listwise item comparisons. Instead, we must infer user preferences by observing utterances during a NL system-user dialogue. At turn $t$ of a dialogue, we let $q^{t}$ and $r^{t}$ be the system and user utterance, respectively, with $\mathbf{q}^{t}=[q^{1},...,q^{t}]$ and $\mathbf{r}^{t}=[r^{1},...,r^{t}]$ representing all system and user utterances up to $t$. In this paper, we call $q^{t}$ the query and $r^{t}$ the response, though extensions to more generic dialogues (e.g., when users can also ask queries) are discussed in Section 7. We let $\mathcal{H}^{t}=(\mathbf{q}^{t},\mathbf{r}^{t})$ be the conversation history at turn $t$.
To formulate NL-PE as a Bayesian optimization problem, we place a prior belief on the user’s utilities $p(\mathbf{u}|\mathbf{x})$, conditioned on item descriptions since they are available before the dialogue begins. We then assume an observation model that gives the likelihood $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t}$), letting us define the posterior utility belief as
(3) $p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t})\propto p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u,\mathbf{q}^{t}})p(\mathbf{u}|\mathbf{x}).$
This posterior informs an acquisition function $\gamma(\mathbf{x},\mathcal{H}^{t})$ which generates [^3] To represent the generative acquisition of NL outputs, we deviate from the conventional definition of acquisition functions as mapping to $\mathbb{R}.$ a new NL query
(4) $q^{t+1}=\gamma(\mathbf{x},\mathcal{H}^{t}),$
to systematically search for $i^{*}$. The preference beliefs also let us define an Expected Utility (EU) $\mu_{i}^{t}$ for every item as
(5) $\mu_{i}^{t}=\mathbb{E}_{p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t})}[u_{i}],$
which allows the top-$k$ items to be recommended at any turn based on their expected utilities.
Our Bayesian optimization NL-PE paradigm lets us formalize several key questions, including:
  1. (1)
    How do we represent beliefs $p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t})$ in user-item utilities $\mathbf{u}$, given NL item descriptions $\mathbf{x}$ and a dialogue $\mathcal{H}^{t}$?
  2. (2)
    What are effective models for the likelihood $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t})$ of observed responses $\mathbf{r}^{t}$ given $\mathbf{x}$, $\mathbf{q}^{t}$, and user utilities $\mathbf{u}$?
  3. (3)
    How can our beliefs inform the generative acquisition of NL queries $q^{t+1}$ given $\mathcal{H}^{t}$ to strategically search for $i^{*}$?
These questions reveal a number of novel research directions discussed further in Section 7. In this paper, we present PEBOL, a NL-PE algorithm based on the above Bayesian optimization NL-PE formalism, and numerically evaluate it against monolithic LLM alternatives through controlled, simulated NL dialogues (cf. Sec. 6).
4. Methodology
Limitations of Monolithic LLM Prompting
An obvious NL-PE approach, described further as baseline in Section 5.1, is to prompt a monolithic LLM with all item descriptions $\mathbf{x}$, dialogue history $\mathcal{H}^{t}$, and instructions to generate a new query at each turn. However, providing all item descriptions $[x_{1},...,x_{N}]$ in the LLM context window is very computationally expensive for all but the smallest item sets. While item knowledge could be internalized through fine-tuning, each item update would imply system retraining. Critically, an LLM’s preference elicitation behaviour cannot be controlled other than by prompt-engineering or further fine-tuning, with neither option offering any guarantees of predictable or interpretable behaviour that balances the exploitation and exploration of user preferences.
PEBOL Overview
We propose to addresses these limitations by augmenting LLM reasoning with a Bayesian Optimization procedure in a novel algorithm, PEBOL, illustrated in Figure 2. At each turn $t$, our algorithm maintains a probabilistic belief state over user preferences as a Beta belief state (cf. Sec. 4.1). This belief state guides an LLM-based acquisition function to generate NL queries explicitly balancing exploration and exploitation to uncover the top user preferences (cf. Sec. 4.2). In addition, our acquisition function reduces the context needed to prompt the LLM in each turn from all $N$ item descriptions $\mathbf{x}$ to a single strategically selected item description $x_{i^{t}}$. PEBOL then uses NLI over elicited NL preferences and item descriptions to map dialogue utterances to numerical observations (c.f. Sec 4.3).
Figure 4. MAP@10 for MonoLLM and PEBOL-P with uncertainty-informed policies (UCB, TS, ER) – all methods show preference learning over time. MonoLLM is outperformed by PEBOL on Yelp and MovieLens, but remains competitive with PEBOL up to the fourth turn on Recipe-MPR until demonstrating a sharp performance drop that takes several turns to recover from.
4.1. Utility Beliefs
4.1.1. Prior Beliefs
Before any dialogue, PEBOL establishes a prior belief $p(\mathbf{u}|\mathbf{x})$ on user-item utilities. We assume item utilities are independent so that
(6) $p(\mathbf{u}|\mathbf{x})=\prod_{i=1}^{N}p(u_{i}|x_{i}),$
and that the prior for each utility $u_{i}$ is a Beta distribution
(7) $p(u_{i}|x_{i})=\text{Beta}(\alpha_{i}^{0},\beta_{i}^{0}).$
Since this paper focuses on fully cold start settings, we assume a uniform Beta prior with $(\alpha_{i}^{0},\beta_{i}^{0})=(1,1)$. Beta distributions, illustrated in Figure 1, lie in the domain $[0,1]$ – a normalized interval for bounded ratings in classical recommendation systems. We can thus interpret utility values of $u_{i}=1$ or $u_{i}=0$ to represent a complete like or dislike of item $i$, respectively, while values $u_{i}\in(0,1)$ provide a strength of preference between these two extremes.
4.1.2. Observation Model
To perform a posterior update on our utility beliefs given observed responses $\mathbf{r}^{t}$, we need an observation model that represents the likelihood $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t})$. Modelling the likelihood of $\mathbf{r}^{t}$ is a challenging task, so we will require some simplifying assumptions. Firstly, we assume that the likelihood of a single response $r^{t}$ is independent from any previous dialogue history $\mathcal{H}^{t-1}$, so that:
(8) $p(\mathbf{r}^{t}|\mathbf{x},\mathbf{u},\mathbf{q}^{t})=\prod_{t^{\prime}=1}^{t}p(r^{t^{\prime}}|\mathbf{x},\mathbf{u},q^{t^{\prime}}).$
Note that this independence assumption will allow incremental posterior belief updates, so that
(9) $p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t})\propto p(r^{t}|\mathbf{x},\mathbf{u},q^{t})p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t-1}).$
4.1.3. Binary Item Response Likelihoods and Posterior Update
With the factorized distributions over item utilities and observational likelihood history now defined, we simply have to provide a concrete observational model of the response likelihood conditioned on the query, item descriptions, and latent utility: $p(r^{t}|\mathbf{x},\mathbf{u},q^{t})$.
[…]
(10) $p(r_{i}^{t}|x_{i},u_{i},q^{t})=\mbox{Bernoulli}(u_{i}).$
With our response likelihood defined, this now leads us to our first pass at a full posterior utility update that we term PEBOL-B for observed Binary rating feedback. Specifically, given observed binary ratings $r^{t}_{i}$, the update at $t=1$ uses the Beta prior (7) with the Bernoulli likelihood (10) to form a standard Beta-Bernoulli conjugate pair and compute the posterior utility belief
(11) $p(u_{i}|x_{i},\mathcal{H}^{1})$ $\propto p(u_{i}|x_{i})p(r^{1}_{i}|x_{i},u_{i},q^{t})$
(12) $=\text{Beta}(\alpha_{i}^{1},\beta_{i}^{1}),$
where $\alpha_{i}^{1}=\alpha_{i}^{0}+r_{i}^{1}$, $\beta_{i}^{1}=\beta_{i}^{0}+(1-r_{i}^{1})$. Subsequent incremental updates updates follow Eq. (9) and use the same conjugacy to give
(13) $p(u_{i}|x_{i},\mathcal{H}^{t})=\text{Beta}(\alpha_{i}^{t},\beta_{i}^{t}),$
where $\alpha_{i}^{t}=\alpha_{i}^{t-1}+r_{i}^{t}$, $\beta_{i}^{t}=\beta_{i}^{t-1}+(1-r_{i}^{t})$.
4.1.4. Natural Language Inference and Probabilistic Posterior Update
As hinted above, effective inference becomes slightly more nuanced since we don’t need to observe an explicit binary response per item in our PEBOL framework. Rather, we receive general preference feedback $r^{t}$ on whether a user generically prefers a text description $q^{t}$ and then leverage an NLI model (Dagan et al., 2005) to infer whether the description $x_{i}$ of item $i$ would be preferred according to this feedback. For instance, for a $(q^{t},r^{t})$ pair (“Want to watch a children’s movie?”,“Yes”), NLI should infer a rating of $r_{1}^{t}=1$ for $x_{1}=$ “The Lion King” and $r_{2}^{t}=0$ for $x_{2}=$ “Titanic”.
To deal with the fact that NLI models actually return an entailment probability, our probabilistic observation variant, PEBOL-P leverages the entailment probability that $r_{i}^{t}=1$ (cf. Sec 4.3), which we denote as $w_{i}^{t}\in[0,1]$. We then simply relax our previous binary observations $r_{i}^{t}$ to continuous $w_{i}^{t}$, thus leading to soft posterior updates $\alpha_{i}^{t}=\alpha_{i}^{t-1}+w_{i}^{t}$, $\beta_{i}^{t}=\beta_{i}^{t-1}+(1-w_{i}^{t})$.
To visually illustrate how this posterior inference process works in practice, Figure 1 shows the effect of PEBOL’s posterior utility belief updates based on NLI for three query-response pairs – we can see the system gaining statistical knowledge about useful items for the user from the dialogue.
Figure 5. MAP@10 for PEBOL-P with various context acquisition policies.
4.2. LLM-Based Acquisition Functions
Recall from Sec. 2.1 that in Bayesian optimization, the posterior informs an acquisition function which determines where to make the next observation. PEBOL generates a new query $q^{t}$ with a two-step acquisition function $\gamma$, first using Bayesian Optimization policies (step 1) based on the posterior utility beliefs $p(\mathbf{u}|\mathbf{x},\mathcal{H}^{t})$ to select NL context, and then using this selected context to guide LLM prompting (step 2). We express the overall acquisition function $\gamma=\gamma^{G}\circ\gamma^{C}$ as a composition of a context acquisition function $\gamma^{C}$ (cf. Sec. 4.2.1) and a NL generation function $\gamma^{G}$ (cf. Sec. 4.2.2).
4.2.1. Context Acquisition via Bayesian Optimization Policies
First, PEBOL harnesses Bayesian optimization policies to select an item description $x_{i^{t}}$ which will be used to prompt an LLM to generate a query about an aspect described by $x_{i^{t}}$ (cf. Sec. 4.2.2). Selecting an item $i^{t}$ whose utility $u_{i^{t}}$ is expected to be near the maximum, $u_{i^{*}}$, will generate exploitation queries asking about properties of items that are likely to be preferred by the user. In contrast, selecting an item $i^{t}$ associated with high uncertainty in its utility $u_{i}^{t}$ will generate exploration queries that probe into properties of items for which user preferences are less known. Thus, strategically selecting $x_{i^{t}}$ allows PEBOL to balance the exploration and exploitation behaviour of NL queries, decreasing the risks of becoming stuck in local optima (over-exploitation) or wasting resources exploring low utility item preferences (over-exploration). We define the item selected by the context acquisition function as
(14) $i^{t}=\gamma^{C}(\mathbf{x},\mathcal{H}^{t}),$
and list several alternatives for $\gamma^{C}$, including the well-known strategies of TS and UCB (Shahriari et al., 2016):
  1. (1)
    Thompson Sampling (TS): First, a sample of each item’s utility $\hat{u}^{t}_{i}$ is taken from the posterior, $\hat{u}^{t}_{i}\sim p(u_{i}|x_{i},\mathcal{H}^{t}).$ Then, the item with the highest sampled utility is selected:
    (15) $i^{t}=\arg\max_{i}\hat{u}^{t}_{i}.$
    TS explores more when beliefs have higher uncertainty and exploits more as the system becomes more confident.
  2. (2)
    Upper Confidence Bound (UCB): Let $P_{k}(\alpha,\beta)$ represent the $k$’th percentile of $\text{Beta}(\alpha,\beta)$, which provides a confidence bound on the posterior. UCB selects the item with the highest confidence bound
    (16) $i^{t}=\arg\max_{i}P_{k}(p(u_{i}|x_{i},\mathcal{H}^{t})),$
    following a balanced strategy because confidence bounds are increased by both high utility and high uncertainty.
  3. (3)
    Entropy Reduction (ER): An explore-only strategy that selects the item with the most uncertain utility:
    (17) $i^{t}=\arg\max_{i}\text{Var}(p(u_{i}|x_{i},\mathcal{H}^{t})).$
  4. (4)
    Greedy: An exploit-only strategy that selects the item with the highest expected utility $\mu_{i}^{t}$ (Eq. 5):
    (18) $i^{t}=\arg\max_{i}\mu_{i}^{t}.$
  5. (5)
    Random: An explore-only heuristic that selects the next item randomly.
4.2.2. Generating Short, Aspect-Based NL Queries
Next, PEBOL prompts an LLM to generate a NL query $q^{t}$ based on the selected item description $x_{i^{t}}$ while also using the dialogue history $\mathcal{H}^{t}$ to avoid repetitive queries. We choose to generate “yes-or-no” queries asking if a user prefers items with some aspect $a^{t}$, which is a short text span extracted dynamically from $x_{i^{t}}$ to be different from any previously queried aspects $a^{1},...,a^{t-1}$. We adopt this query generation strategy to: 1) reduce cognitive load on the user, who may be frustrated by long and specific queries about unfamiliar items and 2) better facilitate NLI through brief, general phrases (Yin et al., 2019). Letting $\phi$ represent the query generation prompt, we let
(19) $q^{t},a^{t}=\gamma^{G}(x_{i^{t}},\mathcal{H}^{t},\phi)$
be the LLM generated query and aspect at turn $t$, with prompting details discussed in Section 5.2.2. An example of such a query and aspect (bold) is “Are you interested in movies with patriotic themes?”, generated by PEBOL in our movie recommendation experiments and shown in Figure 2.
4.3. NL Item-Preference Entailment
4.3.1. Preference Descriptions from Query Response Pairs
Next, PEBOL receives a NL user response $r^{t}$, which it must convert to individual item preference observations. Since the LLM is instructed to generate ”yes-or-no” queries $q^{t}$ asking a user if they like aspect $a^{t}$, we assume the user response will be a ”yes” or a ”no”, and create a NL description of the users preference $\rho^{t}$, letting $\rho^{t}=a^{t}$ if $r^{t}=$“yes”, and $\rho^{t}=\text{concat({``not ''}},a^{t})$ if $r^{t}=$ “no”. For example, given a query that asks if the user prefers the aspect “patriotism” in an item, if the user response is “yes”, then the user preference $\rho^{t}$ is “patriotism”, and “not patriotism” otherwise. This approach produces short, general preference descriptions that are well suited for NLI models (Yin et al., 2019).
4.3.2. Inferring Item Ratings from NL Preferences
Given a NL preference $\rho^{t}$, PEBOL must infer whether the user would like an item described by $x_{i}$. Specifically, PEBOL acquires ratings $\mathbf{w}^{t}=[w^{t}_{1},...,w^{t}_{N}]$ (cf. Sec. 4.1.4) by using NLI to predict whether an item description $x_{i}$ entails (i.e., implies) the preference $\rho^{t}$. For example, we expect that an NLI model would predict that $x_{i}=$The Lion King” entails $\rho^{t}=$“animated” while $x_{j}=$“Titanic” does not, inferring that a user who expressed preference $\rho^{t}$ would like item $i$ but not $j$. We use an NLI model $g_{\theta}(x_{i},\rho^{t})$ to predict the probability $w_{i}^{t}$ that $x_{i}$ entails $\rho^{t}$, and return $r_{i}^{t}=\lfloor w_{i}^{t}\rceil$ in the case of binary observations (PEBOL-B) and $w_{i}^{t}$ in the case of probabilistic observations (PEBOL-P).
4.4. The Complete PEBOL System
This concludes the PEBOL specification – the entire process from prior utility belief to the LLM-based acquisition function generation of a query to the posterior utility update is illustrated in Figure 2.
Figure 6. MAP@10 for PEBOL using binary vs. probabilistic entailment scores. PEBOL-P with the best policy (TS on Yelp and MovieLens, UCB on Recipe-MPR) generally outperforms PEBOL-B.
5. Experimental Methods
We numerically evaluate our PEBOL variations through controlled NL-PE dialogue experiments across multiple datasets and response noise levels – comparing against a monolithic GPT-3.5-turbo NL-PE system, which we refer to as MonoLLM. We do not compare against ConvRec methods (Li et al., 2018; Chen et al., 2019; Wang et al., 2022; Yang et al., 2022) because they are not cold-start systems, requiring observed user-item interactions data to drive their recommendation modules. We also do not base our experiments on ConvRec datasets such as ReDIAL (Li et al., 2018), since they are made up of pre-recorded conversation histories and cannot be used to evaluate active, cold-start NL-PE systems.
5.1. MonoLLM Baseline
A major challenge of using MonoLLM for NL-PE is that item descriptions $\mathbf{x}$ either need to be internalized through training or be provided in the context window (cf. Sec. 4). Since we focus on fully cold-start settings, we test the latter approach as a baseline. Specifically, we prompt the LLM to generate new queries to elicit preferences given the full conversation history $\mathcal{H}^{t}$ and $\mathbf{x}$ (all prompts are shown in the Supplementary Materials). We evaluate recommendation performance after each turn by using another prompt to recommend a list of ten item names from $\mathbf{x}$ given $\mathcal{H}^{t}$. Due to context window limits, this MonoLLM approach is only feasible for small item sets with short item descriptions; thus, we have to limit $|\mathcal{I}|$ to 100 for fair comparison to the MonoLLM baseline.
5.2. Simulation Details
We test PEBOL and MonoLLM through NL-PE dialogues with LLM-simulated users whose item preferences are hidden from the system, evaluating recommendation performance over 10 turns of dialogue, and using GPT-3.5-turbo-0613 for all text generation.
5.2.1. User Simulation
For each experiment, we simulate 100 users, each of which likes a single item $i\in\mathcal{I}$. Each user is simulated by GPT-3.5-turbo-0613, which is given item description $x_{i}$ and instructed to provide only “yes” or “no” responses to a query $q^{t}$ as if it was a user who likes item $i$.
5.2.2. PEBOL Query Generation
Given an item description $x_{i}$ and previously generated aspects $(a^{1},...,a^{t-1})$, an LLM is prompted to generate an aspect $a^{t}$ describing the item $i$ that is no more than 3 words long. The LLM is then prompted again to generate a “yes-or-no” query asking if a user prefers $a^{t}$.
5.2.3. NLI
We use the 400M FAIR mNLI [^4] https://huggingface.co/facebook/bart-large-mnli model to predicts logits for entailment, contradiction, and neutral, and divide these logits by an MNLI temperature $T\in\{1,10,100\}$ As per the FAIR guidelines, we pass the temperature-scaled entailment and contradiction scores through a softmax layer and take the entailment probabilities.
5.2.4. User Response Noise
[…]
5.2.5. Omitting Query History Ablation
We test how tracking query history in PEBOL effects performance with an ablation study that removes previously generated aspects $(a^{1},...,a^{t-1})$ from the aspect extraction prompt.
5.3. Datasets
We obtain item descriptions from three real-world datasets: MovieLens25M [^5] https://grouplens.org/datasets/movielens/25m/, Yelp [^6] https://www.yelp.com/dataset, and Recipe-MPR (Zhang et al., 2023) (example item descriptions from each shown in Table 1 in the Supplementary Materials). After the filtering steps below for Yelp and MovieLens, we randomly sample 100 items to create $\mathbf{x}$. For Yelp, we filter restaurant descriptions to be from a single major North American city and to have at least 50 reviews and five or more category labels. For MovieLens, [^7] For all experiments with MovieLens, we use the 16k version of GPT-3.5-turbo-0613, due to MonoLLM requiring extra context length for $\mathbf{x}.$ we filter movies to be in the 10% by rating count with at least 20 tags, and let movie descriptions use the title, genre labels, and 20 most common user-assigned tags.
Figure 7. The effect of including the generated aspect history in the aspect generation prompt. Including the history improves performance, which we hypothesize is due to reducing repeated or uninformative queries.
Figure 8. The effect of user response noise on MAP@10 – error bars are 95% confidence intervals. At all noise levels, MonoLLM is outperformed by all PEBOL-P methods on Yelp and MovieLens and by UCB, Greedy, and ER on Recipe-MPR.
5.4. Research Questions
Our experiments explore the following research questions (RQs):
  • RQ1: How does PEBOL perform against the GPT-3.5 Mono-LLM baseline?
  • RQ2: Does PEBOL perform better with binary or probabilistic observations, and how sensitive is the latter to temperature?
  • RQ3: How do PEBOL and MonoLLM perform under user response noise?
  • RQ4: How do the context selection policies of TS, UCB, ER, Greedy, and Random effect PEBOL performance?
  • RQ5: How much does PEBOL performance depend on access to the query history during query generation?
6. Experimental Results
6.1. RQ1 - PEBOL vs. MonoLLM
Figure 4 shows MAP@10 over 10 dialogue turns for MonoLLM and PEBOL (UCB,TS,ER), with all methods starting near random guessing (reflecting a cold start) and showing clear preference learning over time. [^8] For each policy, we use the MNLI temperature that performed best on the most datasets (see Supplementary Materials) with continous responses. PEBOL achieves a 131% improvement over MonoLLM in MAP@10 after 10 dialogue turns on Yelp, an 88% improvement on MovieLens, and a 55% improvement on RecipeMPR. MonoLLM also displays a large drop in MAP@10 between turns 4 and 5 on Recipe-MPR which it takes an additional five turns to recover from. In contrast, we do not see any comparably large drops in our methods, indicating that the incremental Bayesian belief updating mechanism of PEBOL may make it less prone to catastrophic mistakes.
6.2. RQ2 - Binary vs. Probabilistic Responses
Figure 6 compares PEBOL performance for PEBOL-B versus PEBOL-P. In general, PEBOL-P performed better, likely due to PEBOL-B discarding valuable information from entailment probabilities. Specifically Figure 6 compares PEBOL-B with each policy against PEBOL-P with the best performing policy on each dataset. On MovieLens, we see at least a 34% improvement in MAP from probabilistic entailment over binary entailment starting from turn 2, though we note that TS with binary updates performs comparably to the best probabilistic methods on Yelp.
To identify our best continuous method, we first compare the performance of PEBOL with PEBOL-P using MNLI temperatures 1, 10, and 100 with each policy. We keep the temperature with the highest MAP@10 at turn 10 for the most datasets: for Yelp and MovieLens, this is TS with MNLI temperature 1; for Recipe-MPR, it is UCB with MNLI temperature 10.
6.3. RQ3 - Effect of User Response Noise
Figure 8 shows the impact of user response noise on MAP@10 at turn 10 – PEBOL generally continues to outperform MonoLLM under user response noise. At turn 10 for all noise levels, MonoLLM is the worst performing method on Yelp and MovieLens, and on RecipeMPR it is always outperformed by UCB, Greedy, and ER. Methods typically maintain their relative performance under various noise levels – though an exception is MovieLens where TS is the best method under zero noise and nearly the worst one at 25% noise (though the margin between the best and worst method in this case is small).
6.4. RQ4 - Comparison of Context Acquisition Policies
Figure 5 compares the performance of various PEBOL context acquisition policies – all policies show active preference learning, other than random item selection on RecipeMPR. There is considerable overlap between methods, however for most turns TS does well on Yelp and MovieLens while being beaten by Greedy, ER, and UCB on Recipe-MPR. As expected due to the randomness in sampling, TS performance is correlated with random item selection, while UCB performs quite similarly to greedy.
6.5. RQ5 - Effect of Aspect History in Query Generation
We see an improvement in performance from including a list of previously generated aspects in the aspect generation prompt. In Figure 7, the average improvement across all turns and all datasets for including the aspect history is 50% for TS and 49% for UCB. Including the aspect generation history helps to avoid repeat queries, which gain no information and would frustrate a user.
7. Conclusion and Future Work
This paper presents a novel Bayesian optimization formalization of natural language (NL) preference elicitation (PE) over arbitrary NL item descriptions, as well as introducing and evaluating PEBOL, an algorithm for NL Preference Elicitation with Bayesian Optimization augmented LLMs – and establishing many new research directions. For instance, while PEBOL uses a pointwise query generation strategy that selects one item description at a time for LLM context, future work can explore LLM-based acquisition functions with pairwise and setwise context selection. Such multi-item context selection would enable contrastive query generation that could better discriminate between item preferences. Another direction for future research is the integration of NL-PE methodologies such as PEBOL into conversational recommendation (ConvRec) system architectures (e.g. (Friedman et al., 2023; Wang et al., 2023)), which must balance many tasks including recommendation, explanation, and personalized question answering. PE in ConvRec systems thus implies that algorithms need to elicit preferences given arbitrary pairs of system-user utterances, in addition to queries and responses.